The stock of Music City Inc. is selling for $37.50. The firm recently paid a dividend of $1.10. What is its implied constant growth rate if the market return is 14 percent?

11.07%

11.4%

10.75%

14.0%
Answer #1

Ans 11.07%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
37.50 = 1.10 / (14%- g)
14% - g = 1.10 / 37.50
14% - g = 2.93%
g = 14% - 2.93%
g = 11.07%